You are taking a quiz with 10 multiple choice questions and each questions has 4 possible answers. If you assume that the questions are independent, what is the standard deviation for quiz scores?
Answer by stanbon(75887)

You are taking a quiz with 10 multiple choice questions and each questions has 4 possible answers. If you assume that the questions are independent, what is the standard deviation for quiz scores?
-----------------------
Binomial Problem with n = 10 and p = 1/4
---
standard deviations = sqrt(npq) = sqrt(10*1/4*3/4)
= sqrt(1.3693)
= 1.1702
=============
Cheers,
Stan H.